Solve 1.7t+8-1.62t=.4t-.32+8
How do I clear the fractions or decimals?

Answer by lwsshak3(11628)

Solve 1.7t+8-1.62t=.4t-.32+8
How do I clear the fractions or decimals?
**
1.7t+8-1.62t=.4t-.32+8
you should move decimals at least two points to the right, so, multiply by 100
170t+800-162t=40t-32+800
170t-162t-40t=-32+800-800
-32t=-32
t=1